2. At the power plant, the water for cooling must be removed at a minimum rate of 65 gal/min. A pump is found in storage thatPlease show work, i will rate. Thank you!
Answer #1

First calculate the rotational speed of thepump required and then calculate the diameter of the pump pulley from rotational speed ratio of the pulley and the motor.

Power input to the motor, Pi = VI (PF) P; = (230V) (22.5)(089) P; = 5175W Rotational speed of motor, Num= 9605 km Rotatioad s

Do I 6 G 4-Dm orfen Np= 3031.8750pm wing. De 4 • (3031-875) Dm Np Dp=1260). - D = 1-266 so, itak Yy round-off Pp = 1.25] w

Bump=5175W pay 25-43 Trump = 16.09Nm or T bumb= 16.09 40.7375 Trump = 11.866 lb-ft Tosque on the pump is 11.866 Uft.